Aggarwal Inc. buys on terms of 2/10 net 30, and it always pays on the 30th day. The CFO calculates that the average amount of costly trade credit carried is $375,000. What is the firm's average accounts payable balance? Assume a 365-day year.
Answer #1
Concepts and reason

Accounting: Accounting is a process of recording the transactions, classifying them in a specific manner, and then it is the process of summarizing and analyzing to interpret the results. It is a process of preserving the accounts.

Liabilities: Liabilities are a company’s owing to the outsiders. It is the obligation raised to a company to settle its payments and dues. Liability is generally classified into long-term liability and short-term liability. Long-term liability is the liability to be settled in the succeeding financial years. Short-term liability or current liability is the liability to be settled in the current financial year.

Fundamentals

Accounts payable: Accounts payable is the amount to be paid by a person or company that has purchased goods or received any services during the future period. It is the liability of the company and thus shown under liabilities in the balance sheet.

Accounts receivable: Accounts receivable is the amount to be received by a company that has sold goods or rendered any services during a period. It is the asset of the company and thus shown under assets in the balance sheet.

Discount rate: The rate at which the purchase price is declined to some extent than that of actual price is known as the discount rate. The rate is fixed by the seller, and the buyer can avail the benefit.

Describe the average accounts payable:

The average accounts payable is utilized because the amount payable can vary throughout the year. It is usually a sum of accounts payable. It is expected to be paid off within a year or within one operating cycle.

Determine the firm’s average accounts payable:

AverageAccountsPayable=TradeCredit+(TradeCredit×DiscountRate×NumberofDays)NumberofMonths=$375,000+($375,000×30×210)12=$375,000+($375,000×30×0.2)12\begin{array}{c}\\{\rm{Average Accounts Payable = Trade Credit + }}\frac{{\left( \begin{array}{l}\\{\rm{Trade Credit \times Discount Rate}}\\\\ \times {\rm{ Number of Days}}\\\end{array} \right)}}{{{\rm{Number of Months}}}}\\\\{\rm{ = \$ 375,000 + }}\frac{{\left( {{\rm{\$ 375,000}} \times 30 \times \frac{2}{{10}}} \right)}}{{12}}\\\\ = {\rm{\$ 375,000 + }}\frac{{\left( {{\rm{\$ 375,000}} \times 30 \times 0.2} \right)}}{{12}}\\\\\\\end{array}

=$375,000+($2,250,00012)=$375,000+$187,500=$562,500\begin{array}{c}\\{\rm{ = \$ 375,000 + }}\left( {\frac{{{\rm{\$ 2,250,000}}}}{{{\rm{12}}}}} \right)\\\\{\rm{ = \$ 375,000 + \$ 187,500}}\\\\{\rm{ = \$ 562,500}}\\\end{array}

Therefore, the firm’s average accounts payable is $562,500.

Ans:

The firm’s average accounts payable is $562,500.